OpenStudy (zab505):

How many roots does the polynomial function, y=(x-8)(x+3)^2 have? A. 2 B. 3 C. 4 D. 1

OpenStudy (jdoe0001):

\(\large { y=(x-8)(x+3)^2\implies y=(x-8)(x+3)(x+3) \\ \quad \\ \qquad \begin{cases} 0=(x-8)\implies 8=x\\ 0=(x+3)\implies -3=x\\ 0=(x+3)\implies -3=x \end{cases}\implies x=\{8,-3\} }\)

OpenStudy (jdoe0001):

it has 3 roots, but one of them has a "multiplicity" of 2 that is is there twice so it really has only 2 roots

OpenStudy (zab505):

Thank you!

OpenStudy (jdoe0001):

yw